Evaluate limz→2i⁡(iz^4+3z^2-10i)



Select one:



A. -12-6i




B. 12+6i




C. 12-6i




D. -12+6i


1
Expert's answer
2022-04-05T06:56:55-0400

"\underset{z\rightarrow 2i}{\lim}\left( iz^4+3z^2-10i \right) =i\left( 2i \right) ^4+3\left( 2i \right) ^2-10i=i\cdot 16+3\cdot \left( -4 \right) -10i=-12+6i\\D\,\,is\,\,correct"